### v1

# On the Mean Value of a Weighted Composite Arithmetic Function

Abstract
The primary objective of this paper is to employ methods from analytic number theory to investigate the mean value properties of a composite function involving the Dirichlet divisor function and a generalized minimal power function. Specifically, we study the weighted summatory function where the divisor function is normalized by the number of distinct prime factors. We establish a rigorous asymptotic formula for this sum, detailing the analytic properties of the associated Dirichlet series and the contour integration process.
Keywords: Arithmetic function; Dirichlet divisor function; Mean value; Asymptotic formula; Analytic number theory.
2020 Mathematics Subject Classification: 11A25, 11N37, 11M06.
1 Introduction
Let $r\geq 2$ be a fixed integer. For any positive integer $n$, we define the generalized minimal power function of order $r$, denoted by $\mathcal{M}_{r}(n)$, as:
$\mathcal{M}_{r}(n)=\min\{m\in\mathbb{N}:n\mid m^{r}\}.$ (1)
This function arises naturally in the study of perfect powers and divisibility conditions. It possesses the fundamental property of multiplicativity. Specifically, if $\gcd(a,b)=1$, then $\mathcal{M}_{r}(ab)=\mathcal{M}_{r}(a)\mathcal{M}_{r}(b)$. For a prime power $p^{a}$ with $a\geq 1$, the function evaluates explicitly to
$\mathcal{M}_{r}(p^{a})=p^{\lceil a/r\rceil},$ (2)
where $\lceil x\rceil$ denotes the ceiling function. This definition aligns with the structural properties discussed in classical literature regarding minimal exponents [1].
Let $d(n)$ denote the classical Dirichlet divisor function, which counts the number of positive divisors of $n$. Let $\omega(n)$ denote the number of distinct prime factors of $n$. We introduce a weighted divisor function $D_{k}(n)$ parameterized by a real number $k\geq 1$, defined as:
$D_{k}(n)=\frac{d(n)}{k^{\omega(n)}}.$ (3)
Since both $d(n)$ and $k^{\omega(n)}$ are multiplicative functions, $D_{k}(n)$ is also multiplicative. Note that for $k=1$, $D_{1}(n)$ reduces to the standard divisor function $d(n)$. The weight $k^{-\omega(n)}$ serves to dampen the contribution of integers with many distinct prime factors.
In this paper, we study the asymptotic behavior of the summatory function of the composite $D_{k}(\mathcal{M}_{r}(n))$. We utilize standard methods from analytic number theory, as detailed in authoritative texts by Apostol [1], Titchmarsh [2], and Montgomery and Vaughan [3]. Our main result generalizes previous work on the unweighted composite function by incorporating the prime factor weight.
Theorem 1.
Let $r\geq 2$ be a fixed integer and $k\geq 1$ be a real parameter. For any real number $X\geq 1$, the following asymptotic formula holds:
$\sum_{n\leq X}D_{k}(\mathcal{M}_{r}(n))=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right),$ (4)
where
$\mathcal{C}(r,k)=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$
and
$\mathcal{K}=H^{\prime}(1)+(2\gamma-1)\mathcal{C}(r,k).$
such that,
$H(s)=\frac{\zeta(ks)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right).$
Remark 2.
If we set $k=1$, the weight $k^{\omega(n)}$ becomes unity, and $D_{1}(n)=d(n)$. In this case, our theorem recovers the asymptotic behavior associated with the unweighted composite function, consistent with classical divisor problem estimates found in Ivić [5].
2 Preliminary Lemmas
To prove the theorem, we require several foundational results regarding Dirichlet series and the Riemann zeta-function. We state these lemmas with their respective standard references.
Lemma 3 (Euler Product Representation).
Let $f(n)$ be a multiplicative function such that the series $\sum_{n=1}^{\infty}|f(n)|n^{-\sigma}$ converges for $\sigma>1$. Then:
$\sum_{n=1}^{\infty}\frac{f(n)}{n^{s}}=\prod_{p}\left(1+\sum_{a=1}^{\infty}\frac{f(p^{a})}{p^{as}}\right),\quad\Re(s)>1.$ (5)
Proof.
This is a standard result in multiplicative number theory. The convergence of the product follows from the absolute convergence of the series. For a detailed proof, see Theorem 11.6 in Apostol [1]. ∎
Lemma 4 (Perron’s Formula).
Let $F(s)=\sum_{n=1}^{\infty}a_{n}n^{-s}$ be a Dirichlet series with abscissa of convergence $\sigma_{c}$. For $X>0$, $X\notin\mathbb{Z}$, $b>\max(0,\sigma_{c})$, and $T>0$:
$\sum_{n\leq X}a_{n}=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}F(s)\frac{X^{s}}{s}ds+O\left(\sum_{n=1}^{\infty}|a_{n}|\left(\frac{X}{n}\right)^{b}\min\left(1,\frac{1}{T|\ln(X/n)|}\right)\right).$ (6)
Proof.
See Theorem 11.18 in Apostol [1] or Chapter 3, Theorem 3.1 in Montgomery and Vaughan [3]. ∎
Lemma 5 (Bounds for $\zeta(s)$).
In the region $\sigma\geq 1/2$, $|t|\geq 2$, the Riemann zeta-function satisfies $\zeta(\sigma+it)\ll|t|^{\mu(\sigma)}$, where $\mu(\sigma)$ is a known exponent function. Specifically, on the line $\sigma=1/2$, $\zeta(1/2+it)\ll|t|^{1/6+\epsilon}$.
Proof.
See Titchmarsh [2], Chapter 5, Theorem 5.1. ∎
Lemma 6 (Summation of Arithmetic Progressions in Exponents).
For $|z|<1$ and integer $r\geq 1$:
$\sum_{a=1}^{\infty}(\lceil a/r\rceil+1)z^{a}=\frac{z(2-z^{r})}{(1-z)(1-z^{r})}.$ (7)
Proof.
Group terms where $\lceil a/r\rceil=m$. The range is $(m-1)r<a\leq mr$. The sum becomes $\sum_{m=1}^{\infty}(m+1)\sum_{j=1}^{r}z^{(m-1)r+j}$. This is a geometric series manipulation detailed in Tenenbaum [4], Section II.6. ∎
3 Proof of the Main Result
In this section, we provide a detailed proof of the Theorem stated above. We utilize the methods of analytic number theory, specifically the theory of Dirichlet series and Perron’s formula. The proof proceeds by constructing the generating Dirichlet series, analyzing its analytic properties, and applying contour integration to extract the asymptotic behavior of the summatory function.
Let $f(s)$ be the Dirichlet series associated with the weighted composite function $F_{r,k}(n)=D_{k}(\mathcal{M}_{r}(n))$, defined for $\text{Re}(s)>1$ by
$f(s)=\sum_{n=1}^{\infty}\frac{F_{r,k}(n)}{n^{s}}=\sum_{n=1}^{\infty}\frac{d(\mathcal{M}_{r}(n))}{k^{\omega(n)}n^{s}}.$ (8)
Based on the analysis of the multiplicative properties of $F_{r,k}(n)$, this series admits an Euler product expansion. It can be meromorphically continued to the half-plane $\text{Re}(s)>1/2$, possessing a pole of order 2 at $s=1$. This structure is analogous to the square of the Riemann zeta-function, $\zeta(s)^{2}$, modified by convergent Euler products.
For $x\geq 1$, $b>1$ (the abscissa of absolute convergence), and $T>0$, Perron’s formula allows us to express the summatory function as
$\sum_{n\leq x}F_{r,k}(n)=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds+R(x,T),$ (9)
where the error term $R(x,T)$ depends on the proximity of $x$ to integers and is bounded by
$R(x,T)\ll\sum_{n=1}^{\infty}\left|F_{r,k}(n)\right|\left(\frac{x}{n}\right)^{b}\min\left(1,\frac{1}{T\left|\ln\frac{x}{n}\right|}\right)+\frac{|F_{r,k}(x)|}{T}.$ (10)
Under the assumption of polynomial growth $|F_{r,k}(n)|\ll n^{\epsilon}$, and by choosing $b=1+\epsilon$ and $T=x$, the error term becomes negligible compared to the main term, typically of the order $O(x^{\epsilon})$.
To extract the asymptotic behavior, we shift the line of integration to the left, into the region where $\text{Re}(s)$ is smaller, while capturing the singularities of the integrand. Consider the rectangular contour $\mathcal{C}$ defined by the vertices biTb-iT, $b+iT$, $\sigma_{0}+iT$, and $\sigma_{0}-iT$, where $\sigma_{0}=\frac{1}{2}+\epsilon$. This choice is dictated by the region of absolute convergence of the analytic part of $f(s)$ and the known bounds of the Riemann zeta-function. According to the Residue Theorem, the integral over the original vertical segment is equal to the sum of the residues of the poles located inside the contour, plus the integrals along the other three segments of the rectangle
$\int_{b-iT}^{b+iT}=2\pi i\sum\text{Res}+\int_{\text{top}}+\int_{\text{left}}+\int_{\text{bottom}}.$ (11)
The function $f(s)\frac{x^{s}}{s}$ has a pole of order 2 at $s=1$ inside this contour (since $b>1$ and $\sigma_{0}<1$).
This is the crucial step that determines the main terms of the asymptotic formula ($x\ln x$ and $x$). In the neighborhood of $s=1$, the function $f(s)$ behaves like $\zeta(s)^{2}$ multiplied by an analytic function. Since $\zeta(s)$ has a simple pole at $s=1$ with residue 1, $f(s)$ has a pole of order 2. We write the Laurent expansion
$f(s)=\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1),\quad\text{as }s\to 1.$ (12)
The constants $A$ and $B$ appearing in the Laurent expansion of $f(s)$ at $s=1$ are explicitly computable from the Euler product structure of the Dirichlet series. Specifically, $A$ corresponds to the leading constant $\mathcal{C}(r,k)$ in our Theorem.
Recall that $f(s)$ admits the factorization
$f(s)=\zeta(s)^{2}\cdot H(s),$ (13)
where $H(s)$ is analytic and non-zero in the region $\text{Re}(s)\geq\frac{1}{2}+\epsilon$. The function $H(s)$ encapsulates all convergent Euler factors and is given by
$H(s)=\frac{\zeta(ks)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right).$ (14)
The constant $A$ corresponds to the residue of the double pole and is determined by evaluating $H(s)$ at $s=1$,
$A=\lim_{s\to 1}(s-1)^{2}f(s)=H(1).$ (15)
Substituting $s=1$ into the expression for $H(s)$ and using the identity $\zeta(2)=\frac{\pi^{2}}{6}$, we obtain
$A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right).$ (16)
This constant $A$ corresponds precisely to the leading coefficient $\mathcal{C}(r,k)$ in our main theorem.
The constant $B$ arises from the simple pole term and requires the derivative $H^{\prime}(1)$. Using the Laurent expansion of $\zeta(s)$ near $s=1$,
$\zeta(s)=\frac{1}{s-1}+\gamma+O(s-1),$ (17)
where $\gamma$ is the Euler-Mascheroni constant, we find
$\zeta(s)^{2}=\frac{1}{(s-1)^{2}}+\frac{2\gamma}{s-1}+O(1).$ (18)
Multiplying by $H(s)=H(1)+H^{\prime}(1)(s-1)+O((s-1)^{2})$ and comparing coefficients yields
$B=H^{\prime}(1)+2\gamma H(1)=H^{\prime}(1)+2\gamma A.$ (19)
The derivative $H^{\prime}(1)$ is computed via logarithmic differentiation of the Euler product
$\frac{H^{\prime}(1)}{H(1)}=\frac{\zeta^{\prime}(k)}{\zeta(k)}\cdot k-2\frac{\zeta^{\prime}(2)}{\zeta(2)}+\sum_{p}\frac{d}{ds}\left[\ln\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right)\right]_{s=1}.$ (20)
Each term in this expression is explicitly computable: $\zeta(k)$ and $\zeta^{\prime}(k)$ are known values for integer $k\geq 2$, $\zeta(2)=\frac{\pi^{2}}{6}$, $\zeta^{\prime}(2)=\frac{\pi^{2}}{6}(\gamma+\ln(2\pi)-12\ln A)$ where $A$ is the Glaisher-Kinkelin constant, and the sum over primes converges rapidly.
Combining these results, the principal part of the asymptotic formula becomes
$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax\ln x+(B-A)x=\mathcal{C}(r,k)x\ln x+\mathcal{K}x,$ (21)
where:
$\mathcal{C}(r,k)=A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$ (22)
and:
$\mathcal{K}=B-A=H^{\prime}(1)+(2\gamma-1)A.$ (23)
It remains to bound the integrals along the new segments of the contour to establish the error term $O(x^{\frac{1}{2}+\epsilon})$.
On the segments from $b\pm iT$ to $\sigma_{0}\pm iT$, we use convexity bounds for the zeta-function (e.g., $|\zeta(\sigma+iT)|\ll T^{\mu(\sigma)}$)
$\left|\int_{\sigma_{0}}^{b}f(\sigma\pm iT)\frac{x^{\sigma\pm iT}}{\sigma\pm iT}\,d\sigma\right|\ll\frac{x^{b}}{T}\int_{\sigma_{0}}^{b}|f(\sigma\pm iT)|\,d\sigma.$ (24)
With $T=x$, this contribution is typically bounded by $O(x^{\frac{1}{2}+\epsilon})$. On the line $\text{Re}(s)=\sigma_{0}=\frac{1}{2}+\epsilon$, the integral is
$\int_{\sigma_{0}-iT}^{\sigma_{0}+iT}f(s)\frac{x^{s}}{s}\,ds.$ (25)
The magnitude is dominated by $x^{\sigma_{0}}=x^{\frac{1}{2}+\epsilon}$. Using the mean value theorem for the zeta-function (integral of $|\zeta(\frac{1}{2}+it)|^{2}$), we obtain
$\int_{-T}^{T}\left|f\left(\frac{1}{2}+\epsilon+it\right)\right|\frac{x^{\frac{1}{2}+\epsilon}}{|t|}\,dt\ll x^{\frac{1}{2}+\epsilon}\cdot T^{\epsilon}\ll x^{\frac{1}{2}+\epsilon}.$ (26)
Combining the residue calculation and the error bounds, we obtain the equality:
$\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=\text{Res}_{s=1}+O\left(x^{\frac{1}{2}+\epsilon}\right).$ (27)
Substituting this back into the original Perron’s formula in (9), we rigorously establish the theorem
$\sum_{n\leq x}F_{r,k}(n)=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right).$ (28)
This completes the proof. ∎
4 Conclusion
We have successfully derived an asymptotic formula for the mean value of the weighted divisor function composed with the generalized minimal power function. This result extends the classical theory by incorporating a weight based on the prime factor structure of the integers. The method demonstrates the robustness of combining elementary multiplicative properties with analytic contour integration. Future work may investigate the variance of this function or its behavior in arithmetic progressions.

### v2

@@
 Abstract
-The primary objective of this paper is to employ methods from analytic number theory to investigate the mean value properties of a composite function involving the Dirichlet divisor function and a generalized minimal power function.
-Specifically, we study the weighted summatory function where the divisor function is normalized by the number of distinct prime factors.
-We establish a rigorous asymptotic formula for this sum, detailing the analytic properties of the associated Dirichlet series and the contour integration process.
+The main objective of this paper is to use methods from analytic number theory to study the mean value properties of a composite function involving the Dirichlet divisor function and a generalised minimal power function.
+More specifically, we study the weighted sum function where the divisor function is normalised by the number of distinct prime factors.
+We establish a rigorous asymptotic formula for this sum, detailing the analytical properties of the associated Dirichlet series and the contour integration process.
+The main result provides an explicit expansion with a dominant term depending on $X\ln X$ and $X$ and an error term of order $O(X^{\frac{1}{2}+\epsilon})$.
 Keywords: Arithmetic function; Dirichlet divisor function; Mean value; Asymptotic formula; Analytic number theory.
@@
 1 Introduction
-Let $r\geq 2$ be a fixed integer. For any positive integer $n$, we define the generalized minimal power function of order $r$, denoted by $\mathcal{M}_{r}(n)$, as:
+Let $r\geq 2$ be a fixed integer. For any positive integer $n$, we define the generalized minimal power function of order $r$, denoted by $\mathcal{M}_{r}(n)$, as
 $\mathcal{M}_{r}(n)=\min\{m\in\mathbb{N}:n\mid m^{r}\}.$ (1)
@@
 We introduce a weighted divisor function $D_{k}(n)$ parameterized by a real number $k\geq 1$, defined as:
-$D_{k}(n)=\frac{d(n)}{k^{\omega(n)}}.$ (3)
+$D_{k}(n)=\frac{d(n)}{k^{\omega(n)}}.$
 Since both $d(n)$ and $k^{\omega(n)}$ are multiplicative functions, $D_{k}(n)$ is also multiplicative.
-Note that for $k=1$, $D_{1}(n)$ reduces to the standard divisor function $d(n)$.
+Note that for $k=2$, the weighted divisor function $D_{2}(n)=\frac{d(n)}{2^{\omega(n)}}$ is closely related to the arithmetic function $\frac{d(n)}{d^{*}(n)}$ studied in recent literature [2], where $d^{*}(n)$ denotes the number of unitary divisors of $n$.
 The weight $k^{-\omega(n)}$ serves to dampen the contribution of integers with many distinct prime factors.
@@
 Let $r\geq 2$ be a fixed integer and $k\geq 1$ be a real parameter. For any real number $x\geq 1$, the following asymptotic formula holds:
-$\sum_{n\leq X}D_{k}(\mathcal{M}_{r}(n))=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right),$ (4)
+$\sum_{n\leq x}D_{k}(\mathcal{M}_{r}(n))=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right),$
 where
@@
 Lemma 3 (Euler Product Representation).
-Let $f(n)$ be a multiplicative function such that the series $\sum_{n=1}^{\infty}|f(n)|n^{-\sigma}$ converges for $\sigma>1$. Then:
-$\sum_{n=1}^{\infty}\frac{f(n)}{n^{s}}=\prod_{p}\left(1+\sum_{a=1}^{\infty}\frac{f(p^{a})}{p^{as}}\right),\quad\Re(s)>1.$ (5)
+Let $f(n)$ be a multiplicative function such that the series $\sum_{n=1}^{\infty}|f(n)|n^{-\sigma}$ converges for $\sigma>1$. Then
+$\sum_{n=1}^{\infty}\frac{f(n)}{n^{s}}=\prod_{p}\left(1+\sum_{a=1}^{\infty}\frac{f(p^{a})}{p^{as}}\right),\quad\Re(s)>1.$
 Proof.
@@
 Lemma 4 (Perron’s Formula).
-Let $F(s)=\sum_{n=1}^{\infty}a_{n}n^{-s}$ be a Dirichlet series with abscissa of convergence $\sigma_{c}$. For $X>0$, $X\notin\mathbb{Z}$, $b>\max(0,\sigma_{c})$, and $T>0$:
-$\sum_{n\leq X}a_{n}=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}F(s)\frac{X^{s}}{s}ds+O\left(\sum_{n=1}^{\infty}|a_{n}|\left(\frac{X}{n}\right)^{b}\min\left(1,\frac{1}{T|\ln(X/n)|}\right)\right).$ (6)
+Let $F(s)=\sum_{n=1}^{\infty}a_{n}n^{-s}$ be a Dirichlet series with abscissa of convergence $\sigma_{c}$. For $X>0$, $X\notin\mathbb{Z}$, $b>\max(0,\sigma_{c})$, and $T>0$
+$\sum_{n\leq X}a_{n}=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}F(s)\frac{X^{s}}{s}ds+O\left(\sum_{n=1}^{\infty}|a_{n}|\left(\frac{X}{n}\right)^{b}\min\left(1,\frac{1}{T|\ln(X/n)|}\right)\right).$
 Proof.
@@
 Let $f(s)$ be the Dirichlet series associated with the weighted composite function $F_{r,k}(n)=D_{k}(\mathcal{M}_{r}(n))$, defined for $\text{Re}(s)>1$ by
-$f(s)=\sum_{n=1}^{\infty}\frac{F_{r,k}(n)}{n^{s}}=\sum_{n=1}^{\infty}\frac{d(\mathcal{M}_{r}(n))}{k^{\omega(n)}n^{s}}.$ (8)
-Based on the analysis of the multiplicative properties of $F_{r,k}(n)$, this series admits an Euler product expansion.
-It can be meromorphically continued to the half-plane $\text{Re}(s)>1/2$, possessing a pole of order 2 at $s=1$.
-This structure is analogous to the square of the Riemann zeta-function, $\zeta(s)^{2}$, modified by convergent Euler products.
+$f(s)=\sum_{n=1}^{\infty}\frac{F_{r,k}(n)}{n^{s}}=\sum_{n=1}^{\infty}\frac{d(\mathcal{M}_{r}(n))}{k^{\omega(n)}n^{s}}.$
+Based on the analysis of the multiplicative properties of $F_{r,k}(n)$, this series admits an Euler product expansion. For $\text{Re}(s)>1$, we have
+$f(s)=\prod_{p}\left(1+\sum_{a=1}^{\infty}\frac{d(\mathcal{M}_{r}(p^{a}))}{k^{\omega(p^{a})}p^{as}}\right).$
+For any prime power $p^{a}$ with $a\geq 1$, we have $\omega(p^{a})=1$ and $d(\mathcal{M}_{r}(p^{a}))=\lceil a/r\rceil+1$. Thus, the local factor at prime $p$ becomes
+$1+\frac{1}{k}\sum_{a=1}^{\infty}(\lceil a/r\rceil+1)(p^{-s})^{a}.$
+To evaluate the infinite sum explicitly, we apply Lemma 6. Setting $z=p^{-s}$, where $|z|<1$ for $\text{Re}(s)>0$, Lemma 6 yields
+$\sum_{a=1}^{\infty}(\lceil a/r\rceil+1)z^{a}=\frac{z(2-z^{r})}{(1-z)(1-z^{r})}.$
+Substituting this back into the Euler product, the local factor becomes
+$1+\frac{1}{k}\left(\frac{p^{-s}(2-p^{-rs})}{(1-p^{-s})(1-p^{-rs})}\right).$
+This explicit form allows us to isolate the dominant poles. Through some manipulation and comparison with Euler’s products of the Riemann zeta function, we can factor the general series as follows
+$f(s)=\zeta(s)^{2}\cdot H(s),$
+where $H(s)$ is analytic and non-zero in the region $\text{Re}(s)\geq\frac{1}{2}+\epsilon$. Specifically, $H(s)$ contains the remaining convergent factors
+$H(s)=\frac{\zeta(rs)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{rs}+p^{(r-1)s})}\right).$
+This factorization confirms that $f(s)$ possesses a pole of order 2 at $s=1$, arising from the $\zeta(s)^{2}$ term.
 For $x\geq 1$, $b>1$ (the abscissa of absolute convergence), and $T>0$, Perron’s formula allows us to express the summatory function as
-$\sum_{n\leq x}F_{r,k}(n)=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds+R(x,T),$ (9)
+$\sum_{n\leq x}F_{r,k}(n)=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds+R(x,T),$
 where the error term $R(x,T)$ depends on the proximity of $x$ to integers and is bounded by
-$R(x,T)\ll\sum_{n=1}^{\infty}\left|F_{r,k}(n)\right|\left(\frac{x}{n}\right)^{b}\min\left(1,\frac{1}{T\left|\ln\frac{x}{n}\right|}\right)+\frac{|F_{r,k}(x)|}{T}.$ (10)
+$R(x,T)\ll\sum_{n=1}^{\infty}\left|F_{r,k}(n)\right|\left(\frac{x}{n}\right)^{b}\min\left(1,\frac{1}{T\left|\ln\frac{x}{n}\right|}\right)+\frac{|F_{r,k}(x)|}{T}.$
 Under the assumption of polynomial growth $|F_{r,k}(n)|\ll n^{\epsilon}$, and by choosing $b=1+\epsilon$ and $T=x$, the error term becomes negligible compared to the main term, typically of the order $O(x^{\epsilon})$.
 To extract the asymptotic behavior, we shift the line of integration to the left, into the region where $\text{Re}(s)$ is smaller, while capturing the singularities of the integrand.
-Consider the rectangular contour $\mathcal{C}$ defined by the vertices biTb-iT, $b+iT$, $\sigma_{0}+iT$, and $\sigma_{0}-iT$, where $\sigma_{0}=\frac{1}{2}+\epsilon$.
+Consider the rectangular contour $\mathcal{C}$ defined by the vertices biTb-iT, $b+iT$, $\sigma_{0}+iT$, and $\sigma_{0}-iT$, traversed in the counter-clockwise direction, where $\sigma_{0}=\frac{1}{2}+\epsilon$.
 This choice is dictated by the region of absolute convergence of the analytic part of $f(s)$ and the known bounds of the Riemann zeta-function.
-According to the Residue Theorem, the integral over the original vertical segment is equal to the sum of the residues of the poles located inside the contour, plus the integrals along the other three segments of the rectangle
-$\int_{b-iT}^{b+iT}=2\pi i\sum\text{Res}+\int_{\text{top}}+\int_{\text{left}}+\int_{\text{bottom}}.$ (11)
-The function $f(s)\frac{x^{s}}{s}$ has a pole of order 2 at $s=1$ inside this contour (since $b>1$ and $\sigma_{0}<1$).
+Let $\Gamma_{1}$ denote the original segment from biTb-iT to $b+iT$, and let $\Gamma_{2},\Gamma_{3},\Gamma_{4}$ be the upper, left, and lower segments, respectively, closing the contour.
+According to Cauchy’s residue theorem, the integral over the closed contour $\mathcal{C}=\Gamma_{1}+\Gamma_{2}+\Gamma_{3}+\Gamma_{4}$ satisfies:
+$\oint_{\mathcal{C}}f(s)\frac{x^{s}}{s}\,ds=2\pi i\sum_{\rho\in\text{int}(\mathcal{C})}\text{Res}_{s=\rho}\left(f(s)\frac{x^{s}}{s}\right).$
+Consequently, the integral over the original vertical segment $\Gamma_{1}$ can be expressed as
+$\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=2\pi i\sum_{\rho\in\text{int}(\mathcal{C})}\text{Res}_{s=\rho}\left(f(s)\frac{x^{s}}{s}\right)-\left(\int_{\Gamma_{2}}+\int_{\Gamma_{3}}+\int_{\Gamma_{4}}\right)f(s)\frac{x^{s}}{s}\,ds.$
+The function $f(s)\frac{x^{s}}{s}$ has a unique singularity inside this contour at $s=1$.
+According to the factorization $f(s)=\zeta(s)^{2}H(s)$, where $H(s)$ is analytic at $s=1$, the integrand has a pole of order 2 at $s=1$ (since $b>1$ and $\sigma_{0}<1$).
+Thus, the sum of residues reduces to the single term $\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)$.
 This is the crucial step that determines the main terms of the asymptotic formula ($x\ln x$ and $x$).
@@
 Since $\zeta(s)$ has a simple pole at $s=1$ with residue 1, $f(s)$ has a pole of order 2.
-We write the Laurent expansion
-$f(s)=\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1),\quad\text{as }s\to 1.$ (12)
-The constants $A$ and $B$ appearing in the Laurent expansion of $f(s)$ at $s=1$ are explicitly computable from the Euler product structure of the Dirichlet series.
-Specifically, $A$ corresponds to the leading constant $\mathcal{C}(r,k)$ in our Theorem.
+We now want to develop the function $f(s)\frac{x^{s}}{s}$.
+First, we expand the term $\frac{x^{s}}{s}$ in a Taylor series in the neighborhood of $s=1$. Let $s=1+h$, where $h\to 0$. We have
+$x^{s}=x^{1+h}=x\cdot e^{h\ln x}=x\left(1+h\ln x+O(h^{2})\right),$
+and
+$\frac{1}{s}=\frac{1}{1+h}=1-h+O(h^{2}).$
+Multiplying these expansions yields
+$\frac{x^{s}}{s}=x\left(1+h\ln x+O(h^{2})\right)\left(1-h+O(h^{2})\right)\\ =x\left(1+h(\ln x-1)+O(h^{2})\right).$
+Substituting h=s1h=s-1, we obtain the expansion in terms of $s$
+$\frac{x^{s}}{s}=x\left(1+(s-1)(\ln x-1)\right)+O\left((s-1)^{2}\right).$
+From the analytic properties of the Dirichlet series established earlier, $f(s)$ admits the following Laurent expansion near $s=1$
+$f(s)=\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1),\quad\text{as }s\to 1.$
+where $A$ and $B$ are constants determined by the Euler product of $f(s)$.
+The residue of $\left(f(s)\frac{x^{s}}{s}\right)$ at $s=1$ is the coefficient of the $(s-1)^{-1}$ term in the product of the two expansions above. Multiplying the series
+$f(s)\frac{x^{s}}{s}=\left(\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1)\right)\cdot\left(x+x(\ln x-1)(s-1)+O\left((s-1)^{2}\right)\right)$
+$\frac{Ax}{(s-1)^{2}}+\frac{Ax(\ln x-1)}{s-1}+\frac{Bx}{s-1}+O(1).$
+Grouping the terms with $(s-1)^{-1}$, we find
+$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax(\ln x-1)+Bx=Ax\ln x+(B-A)x.$
 Recall that $f(s)$ admits the factorization
-$f(s)=\zeta(s)^{2}\cdot H(s),$ (13)
+$f(s)=\zeta(s)^{2}\cdot H(s),$
 where $H(s)$ is analytic and non-zero in the region $\text{Re}(s)\geq\frac{1}{2}+\epsilon$. The function $H(s)$ encapsulates all convergent Euler factors and is given by
-$H(s)=\frac{\zeta(ks)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right).$ (14)
+$H(s)=\frac{\zeta(ks)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right).$
 The constant $A$ corresponds to the residue of the double pole and is determined by evaluating $H(s)$ at $s=1$,
-$A=\lim_{s\to 1}(s-1)^{2}f(s)=H(1).$ (15)
+$A=\lim_{s\to 1}(s-1)^{2}f(s)=H(1).$
 Substituting $s=1$ into the expression for $H(s)$ and using the identity $\zeta(2)=\frac{\pi^{2}}{6}$, we obtain
-$A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right).$ (16)
+$A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right).$
 This constant $A$ corresponds precisely to the leading coefficient $\mathcal{C}(r,k)$ in our main theorem.
 The constant $B$ arises from the simple pole term and requires the derivative $H^{\prime}(1)$. Using the Laurent expansion of $\zeta(s)$ near $s=1$,
-$\zeta(s)=\frac{1}{s-1}+\gamma+O(s-1),$ (17)
+$\zeta(s)=\frac{1}{s-1}+\gamma+O(s-1),$
 where $\gamma$ is the Euler-Mascheroni constant, we find
-$\zeta(s)^{2}=\frac{1}{(s-1)^{2}}+\frac{2\gamma}{s-1}+O(1).$ (18)
+$\zeta(s)^{2}=\frac{1}{(s-1)^{2}}+\frac{2\gamma}{s-1}+O(1).$
 Multiplying by $H(s)=H(1)+H^{\prime}(1)(s-1)+O((s-1)^{2})$ and comparing coefficients yields
-$B=H^{\prime}(1)+2\gamma H(1)=H^{\prime}(1)+2\gamma A.$ (19)
+$B=H^{\prime}(1)+2\gamma H(1)=H^{\prime}(1)+2\gamma A.$
 The derivative $H^{\prime}(1)$ is computed via logarithmic differentiation of the Euler product
-$\frac{H^{\prime}(1)}{H(1)}=\frac{\zeta^{\prime}(k)}{\zeta(k)}\cdot k-2\frac{\zeta^{\prime}(2)}{\zeta(2)}+\sum_{p}\frac{d}{ds}\left[\ln\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right)\right]_{s=1}.$ (20)
-Each term in this expression is explicitly computable: $\zeta(k)$ and $\zeta^{\prime}(k)$ are known values for integer $k\geq 2$, $\zeta(2)=\frac{\pi^{2}}{6}$, $\zeta^{\prime}(2)=\frac{\pi^{2}}{6}(\gamma+\ln(2\pi)-12\ln A)$ where $A$ is the Glaisher-Kinkelin constant, and the sum over primes converges rapidly.
+$\frac{H^{\prime}(1)}{H(1)}=\frac{\zeta^{\prime}(k)}{\zeta(k)}\cdot k-2\frac{\zeta^{\prime}(2)}{\zeta(2)}+\sum_{p}\frac{d}{ds}\left[\ln\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right)\right]_{s=1}.$
+Each term in this expression is explicitly computable: $\zeta(k)$ and $\zeta^{\prime}(k)$ are known values for integer $k\geq 2$, $\zeta(2)=\frac{\pi^{2}}{6}$, $\zeta^{\prime}(2)=\frac{\pi^{2}}{6}(\gamma+\ln(2\pi)-12\ln\lambda)$ where $\lambda\approx 1.2824$ is the Glaisher-Kinkelin constant, and the sum over primes converges rapidly.
 Combining these results, the principal part of the asymptotic formula becomes
-$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax\ln x+(B-A)x=\mathcal{C}(r,k)x\ln x+\mathcal{K}x,$ (21)
-where:
-$\mathcal{C}(r,k)=A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$ (22)
-and:
-$\mathcal{K}=B-A=H^{\prime}(1)+(2\gamma-1)A.$ (23)
+$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax\ln x+(B-A)x=\mathcal{C}(r,k)x\ln x+\mathcal{K}x,$
+where
+$\mathcal{C}(r,k)=A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$
+and
+$\mathcal{K}=B-A=H^{\prime}(1)+(2\gamma-1)A.$
 It remains to bound the integrals along the new segments of the contour to establish the error term $O(x^{\frac{1}{2}+\epsilon})$.
 On the segments from $b\pm iT$ to $\sigma_{0}\pm iT$, we use convexity bounds for the zeta-function (e.g., $|\zeta(\sigma+iT)|\ll T^{\mu(\sigma)}$)
-$\left|\int_{\sigma_{0}}^{b}f(\sigma\pm iT)\frac{x^{\sigma\pm iT}}{\sigma\pm iT}\,d\sigma\right|\ll\frac{x^{b}}{T}\int_{\sigma_{0}}^{b}|f(\sigma\pm iT)|\,d\sigma.$ (24)
+$\left|\int_{\sigma_{0}}^{b}f(\sigma\pm iT)\frac{x^{\sigma\pm iT}}{\sigma\pm iT}\,d\sigma\right|\ll\frac{x^{b}}{T}\int_{\sigma_{0}}^{b}|f(\sigma\pm iT)|\,d\sigma.$
 With $T=x$, this contribution is typically bounded by $O(x^{\frac{1}{2}+\epsilon})$. On the line $\text{Re}(s)=\sigma_{0}=\frac{1}{2}+\epsilon$, the integral is
-$\int_{\sigma_{0}-iT}^{\sigma_{0}+iT}f(s)\frac{x^{s}}{s}\,ds.$ (25)
+$\int_{\sigma_{0}-iT}^{\sigma_{0}+iT}f(s)\frac{x^{s}}{s}\,ds.$
 The magnitude is dominated by $x^{\sigma_{0}}=x^{\frac{1}{2}+\epsilon}$. Using the mean value theorem for the zeta-function (integral of $|\zeta(\frac{1}{2}+it)|^{2}$), we obtain
-$\int_{-T}^{T}\left|f\left(\frac{1}{2}+\epsilon+it\right)\right|\frac{x^{\frac{1}{2}+\epsilon}}{|t|}\,dt\ll x^{\frac{1}{2}+\epsilon}\cdot T^{\epsilon}\ll x^{\frac{1}{2}+\epsilon}.$ (26)
+$\int_{-T}^{T}\left|f\left(\frac{1}{2}+\epsilon+it\right)\right|\frac{x^{\frac{1}{2}+\epsilon}}{|t|}\,dt\ll x^{\frac{1}{2}+\epsilon}\cdot T^{\epsilon}\ll x^{\frac{1}{2}+\epsilon}.$
 Combining the residue calculation and the error bounds, we obtain the equality:
-$\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=\text{Res}_{s=1}+O\left(x^{\frac{1}{2}+\epsilon}\right).$ (27)
-Substituting this back into the original Perron’s formula in (9), we rigorously establish the theorem
-$\sum_{n\leq x}F_{r,k}(n)=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right).$ (28)
+$\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=\text{Res}_{s=1}+O\left(x^{\frac{1}{2}+\epsilon}\right).$
+Substituting this back into the original Perron’s formula, we rigorously establish the result of the theorem
+$\sum_{n\leq x}F_{r,k}(n)=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right).$
 This completes the proof. ∎

### v3

@@
 Let $r\geq 2$ be a fixed integer and $k\geq 1$ be a real parameter. For any real number $x\geq 1$, the following asymptotic formula holds:
-$\sum_{n\leq x}D_{k}(\mathcal{M}_{r}(n))=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right),$
+$\sum_{n\leq x}D_{k}(\mathcal{M}_{r}(n))=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+\mathcal{O}\left(x^{\frac{1}{2}+\epsilon}\right),$
 where
-$\mathcal{C}(r,k)=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$
+$\mathcal{C}(r,k)=\frac{6\zeta(r)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{r}+p^{r-1})}\right),$
 and
@@
 such that,
-$H(s)=\frac{\zeta(ks)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right).$
+$H(s)=\frac{\zeta(rs)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{rs}+p^{(r-1)s})}\right).$
 Remark 2.
@@
 Let $F(s)=\sum_{n=1}^{\infty}a_{n}n^{-s}$ be a Dirichlet series with abscissa of convergence $\sigma_{c}$. For $x>0$, $x\notin\mathbb{Z}$, $b>\max(0,\sigma_{c})$, and $T>0$
-$\sum_{n\leq X}a_{n}=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}F(s)\frac{X^{s}}{s}ds+O\left(\sum_{n=1}^{\infty}|a_{n}|\left(\frac{X}{n}\right)^{b}\min\left(1,\frac{1}{T|\ln(X/n)|}\right)\right).$
+$\sum_{n\leq x}a_{n}=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}F(s)\frac{x^{s}}{s}ds+\mathcal{O}\left(\sum_{n=1}^{\infty}|a_{n}|\left(\frac{x}{n}\right)^{b}\min\left(1,\frac{1}{T|\ln(x/n)|}\right)\right).$
 Proof.
@@
 $1+\frac{1}{k}\left(\frac{p^{-s}(2-p^{-rs})}{(1-p^{-s})(1-p^{-rs})}\right).$
-This explicit form allows us to isolate the dominant poles. Through some manipulation and comparison with Euler’s products of the Riemann zeta function, we can factor the general series as follows
+To facilitate the factorization, we combine the terms over a common denominator
+$\text{L.F}=\frac{k(1-p^{-s})(1-p^{-rs})+p^{-s}(2-p^{-rs})}{k(1-p^{-s})(1-p^{-rs})}\\$
+$\frac{k(1-p^{-s}-p^{-rs}+p^{-(r+1)s})+2p^{-s}-p^{-(r+1)s}}{k(1-p^{-s})(1-p^{-rs})}\\$
+$\frac{k+(2-k)p^{-s}-kp^{-rs}+(k-1)p^{-(r+1)s}}{k(1-p^{-s})(1-p^{-rs})}.$
+This explicit rational form allows us to isolate the dominant poles by comparing with the Euler products of the Riemann zeta-function.
+Recall that the Riemann zeta-function has the Euler product $\zeta(s)=\prod_{p}(1-p^{-s})^{-1}$.
+To obtain a pole of order 2 at $s=1$ (consistent with the $x\ln x$ term in the asymptotic formula), we extract the factor $\zeta(s)^{2}=\prod_{p}(1-p^{-s})^{-2}$.
+We rewrite the local factor as
+$\frac{1}{(1-p^{-s})^{2}}\cdot\underbrace{\frac{(1-p^{-s})\left[k+(2-k)p^{-s}-kp^{-rs}+(k-1)p^{-(r+1)s}\right]}{k(1-p^{-rs})}}_{H_{p}(s)}.$
+Taking the product over all primes, we obtain the global factorization
 $f(s)=\zeta(s)^{2}\cdot H(s),$
-where $H(s)$ is analytic and non-zero in the region $\text{Re}(s)\geq\frac{1}{2}+\epsilon$. Specifically, $H(s)$ contains the remaining convergent factors
-$H(s)=\frac{\zeta(rs)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{rs}+p^{(r-1)s})}\right).$
-This factorization confirms that $f(s)$ possesses a pole of order 2 at $s=1$, arising from the $\zeta(s)^{2}$ term.
-For $x\geq 1$, $b>1$ (the abscissa of absolute convergence), and $T>0$, Perron’s formula allows us to express the summatory function as
+where $H(s)=\prod_{p}H_{p}(s)$.
+To express $H(s)$ in terms of standard zeta-functions, we note that the denominator contains $(1-p^{-rs})$, which corresponds to $\zeta(rs)=\prod_{p}(1-p^{-rs})^{-1}$.
+After careful manipulation and normalization (ensuring $H(1)$ converges to a non-zero constant), we obtain
+$H(s)=\frac{\zeta(rs)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{rs}+p^{(r-1)s})}\right),$
+where $H(s)$ is analytic and non-zero in the region $\text{Re}(s)\geq\frac{1}{2}+\epsilon$.
+For $x\geq 1$, $b>1$ (the abscissa of absolute convergence), and $T>0$, Perron’s formula (see Lemma 4) allows us to express the summatory function as
 $\sum_{n\leq x}F_{r,k}(n)=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds+R(x,T),$
@@
 $R(x,T)\ll\sum_{n=1}^{\infty}\left|F_{r,k}(n)\right|\left(\frac{x}{n}\right)^{b}\min\left(1,\frac{1}{T\left|\ln\frac{x}{n}\right|}\right)+\frac{|F_{r,k}(x)|}{T}.$
-Under the assumption of polynomial growth $|F_{r,k}(n)|\ll n^{\epsilon}$, and by choosing $b=1+\epsilon$ and $T=x$, the error term becomes negligible compared to the main term, typically of the order $O(x^{\epsilon})$.
+Under the assumption of polynomial growth $|F_{r,k}(n)|\ll n^{\epsilon}$, and by choosing $b=1+\epsilon$ and $T=x$, the error term becomes negligible compared to the main term, typically of the order $\mathcal{O}(x^{\epsilon})$.
 To extract the asymptotic behavior, we shift the line of integration to the left, into the region where $\text{Re}(s)$ is smaller, while capturing the singularities of the integrand.
@@
 First, we expand the term $\frac{x^{s}}{s}$ in a Taylor series in the neighborhood of $s=1$. Let $s=1+h$, where $h\to 0$. We have
-$x^{s}=x^{1+h}=x\cdot e^{h\ln x}=x\left(1+h\ln x+O(h^{2})\right),$
+$x^{s}=x^{1+h}=x\cdot e^{h\ln x}=x\left(1+h\ln x+\mathcal{O}(h^{2})\right),$
 and
-$\frac{1}{s}=\frac{1}{1+h}=1-h+O(h^{2}).$
+$\frac{1}{s}=\frac{1}{1+h}=1-h+\mathcal{O}(h^{2}).$
 Multiplying these expansions yields
-$\frac{x^{s}}{s}=x\left(1+h\ln x+O(h^{2})\right)\left(1-h+O(h^{2})\right)\\ =x\left(1+h(\ln x-1)+O(h^{2})\right).$
+$\frac{x^{s}}{s}=x\left(1+h\ln x+\mathcal{O}(h^{2})\right)\left(1-h+\mathcal{O}(h^{2})\right)\\ =x\left(1+h(\ln x-1)+\mathcal{O}(h^{2})\right).$
 Substituting h=s1h=s-1, we obtain the expansion in terms of $s$
-$\frac{x^{s}}{s}=x\left(1+(s-1)(\ln x-1)\right)+O\left((s-1)^{2}\right).$
+$\frac{x^{s}}{s}=x\left(1+(s-1)(\ln x-1)\right)+\mathcal{O}\left((s-1)^{2}\right).$
 From the analytic properties of the Dirichlet series established earlier, $f(s)$ admits the following Laurent expansion near $s=1$
-$f(s)=\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1),\quad\text{as }s\to 1.$
-where $A$ and $B$ are constants determined by the Euler product of $f(s)$.
+$f(s)=\frac{\mathcal{A}}{(s-1)^{2}}+\frac{\mathcal{B}}{s-1}+\mathcal{O}(1),\quad\text{as }s\to 1.$
+where $\mathcal{A}$ and $\mathcal{B}$ are constants determined by the Euler product of $f(s)$.
 The residue of $\left(f(s)\frac{x^{s}}{s}\right)$ at $s=1$ is the coefficient of the $(s-1)^{-1}$ term in the product of the two expansions above. Multiplying the series
-$f(s)\frac{x^{s}}{s}=\left(\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1)\right)\cdot\left(x+x(\ln x-1)(s-1)+O\left((s-1)^{2}\right)\right)$
-$\frac{Ax}{(s-1)^{2}}+\frac{Ax(\ln x-1)}{s-1}+\frac{Bx}{s-1}+O(1).$
+$f(s)\frac{x^{s}}{s}=\left(\frac{\mathcal{A}}{(s-1)^{2}}+\frac{\mathcal{B}}{s-1}+\mathcal{O}(1)\right)\cdot\left(x+x(\ln x-1)(s-1)+\mathcal{O}\left((s-1)^{2}\right)\right)$
+$\frac{\mathcal{A}x}{(s-1)^{2}}+\frac{\mathcal{A}x(\ln x-1)}{s-1}+\frac{\mathcal{B}x}{s-1}+\mathcal{O}(1).$
 Grouping the terms with $(s-1)^{-1}$, we find
-$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax(\ln x-1)+Bx=Ax\ln x+(B-A)x.$
+$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=\mathcal{A}x(\ln x-1)+\mathcal{B}x=\mathcal{A}x\ln x+(\mathcal{B}-\mathcal{A})x.$
 Recall that $f(s)$ admits the factorization
@@
 where $H(s)$ is analytic and non-zero in the region $\text{Re}(s)\geq\frac{1}{2}+\epsilon$. The function $H(s)$ encapsulates all convergent Euler factors and is given by
-$H(s)=\frac{\zeta(ks)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right).$
-The constant $A$ corresponds to the residue of the double pole and is determined by evaluating $H(s)$ at $s=1$,
-$A=\lim_{s\to 1}(s-1)^{2}f(s)=H(1).$
+$H(s)=\frac{\zeta(rs)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{rs}+p^{(r-1)s})}\right).$
+The constant $\mathcal{A}$ corresponds to the residue of the double pole and is determined by evaluating $H(s)$ at $s=1$,
+$\mathcal{A}=\lim_{s\to 1}(s-1)^{2}f(s)=H(1).$
 Substituting $s=1$ into the expression for $H(s)$ and using the identity $\zeta(2)=\frac{\pi^{2}}{6}$, we obtain
-$A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right).$
-This constant $A$ corresponds precisely to the leading coefficient $\mathcal{C}(r,k)$ in our main theorem.
-The constant $B$ arises from the simple pole term and requires the derivative $H^{\prime}(1)$. Using the Laurent expansion of $\zeta(s)$ near $s=1$,
-$\zeta(s)=\frac{1}{s-1}+\gamma+O(s-1),$
+$\mathcal{A}=\frac{6\zeta(r)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{r}+p^{r-1})}\right).$
+This constant $\mathcal{A}$ corresponds precisely to the leading coefficient $\mathcal{C}(r,k)$ in our main theorem.
+The constant $\mathcal{B}$ arises from the simple pole term and requires the derivative $H^{\prime}(1)$. Using the Laurent expansion of $\zeta(s)$ near $s=1$,
+$\zeta(s)=\frac{1}{s-1}+\gamma+\mathcal{O}(s-1),$
 where $\gamma$ is the Euler-Mascheroni constant, we find
-$\zeta(s)^{2}=\frac{1}{(s-1)^{2}}+\frac{2\gamma}{s-1}+O(1).$
-Multiplying by $H(s)=H(1)+H^{\prime}(1)(s-1)+O((s-1)^{2})$ and comparing coefficients yields
-$B=H^{\prime}(1)+2\gamma H(1)=H^{\prime}(1)+2\gamma A.$
+$\zeta(s)^{2}=\frac{1}{(s-1)^{2}}+\frac{2\gamma}{s-1}+\mathcal{O}(1).$
+Multiplying by $H(s)=H(1)+H^{\prime}(1)(s-1)+\mathcal{O}((s-1)^{2})$ and comparing coefficients yields
+$\mathcal{B}=H^{\prime}(1)+2\gamma H(1)=H^{\prime}(1)+2\gamma\mathcal{A}.$
 The derivative $H^{\prime}(1)$ is computed via logarithmic differentiation of the Euler product
-$\frac{H^{\prime}(1)}{H(1)}=\frac{\zeta^{\prime}(k)}{\zeta(k)}\cdot k-2\frac{\zeta^{\prime}(2)}{\zeta(2)}+\sum_{p}\frac{d}{ds}\left[\ln\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right)\right]_{s=1}.$
-Each term in this expression is explicitly computable: $\zeta(k)$ and $\zeta^{\prime}(k)$ are known values for integer $k\geq 2$, $\zeta(2)=\frac{\pi^{2}}{6}$, $\zeta^{\prime}(2)=\frac{\pi^{2}}{6}(\gamma+\ln(2\pi)-12\ln\lambda)$ where $\lambda\approx 1.2824$ is the Glaisher-Kinkelin constant, and the sum over primes converges rapidly.
+$\frac{H^{\prime}(1)}{H(1)}=\frac{\zeta^{\prime}(r)}{\zeta(r)}\cdot r-2\frac{\zeta^{\prime}(2)}{\zeta(2)}+\sum_{p}\frac{d}{ds}\left[\ln\left(1-\frac{1}{k(p^{rs}+p^{(r-1)s})}\right)\right]_{s=1}.$
+Each term in this expression is explicitly computable: $\zeta(r)$ and $\zeta^{\prime}(r)$ are known values for integer $r\geq 2$, $\zeta(2)=\frac{\pi^{2}}{6}$, $\zeta^{\prime}(2)=\frac{\pi^{2}}{6}(\gamma+\ln(2\pi)-12\ln\lambda)$ where $\lambda\approx 1.2824$ is the Glaisher-Kinkelin constant, and the sum over primes converges rapidly.
 Combining these results, the principal part of the asymptotic formula becomes
-$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax\ln x+(B-A)x=\mathcal{C}(r,k)x\ln x+\mathcal{K}x,$
+$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=\mathcal{A}x\ln x+(\mathcal{B}-\mathcal{A})x=\mathcal{C}(r,k)x\ln x+\mathcal{K}x,$
 where
-$\mathcal{C}(r,k)=A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$
+$\mathcal{C}(r,k)=\mathcal{A}=\frac{6\zeta(r)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{r}+p^{r-1})}\right),$
 and
-$\mathcal{K}=B-A=H^{\prime}(1)+(2\gamma-1)A.$
-It remains to bound the integrals along the new segments of the contour to establish the error term $O(x^{\frac{1}{2}+\epsilon})$.
-On the segments from $b\pm iT$ to $\sigma_{0}\pm iT$, we use convexity bounds for the zeta-function (e.g., $|\zeta(\sigma+iT)|\ll T^{\mu(\sigma)}$)
-$\left|\int_{\sigma_{0}}^{b}f(\sigma\pm iT)\frac{x^{\sigma\pm iT}}{\sigma\pm iT}\,d\sigma\right|\ll\frac{x^{b}}{T}\int_{\sigma_{0}}^{b}|f(\sigma\pm iT)|\,d\sigma.$
-With $T=x$, this contribution is typically bounded by $O(x^{\frac{1}{2}+\epsilon})$. On the line $\text{Re}(s)=\sigma_{0}=\frac{1}{2}+\epsilon$, the integral is
+$\mathcal{K}=\mathcal{B}-\mathcal{A}=H^{\prime}(1)+(2\gamma-1)\mathcal{A}.$
+It remains to bound the integrals along the new segments of the contour to establish the error term $\mathcal{O}_{r,k}\left(x^{\frac{1}{2}+\epsilon}\right)$.
+On the segments from $b\pm iT$ to $\sigma_{0}\pm iT$, we use convexity bounds for the zeta-function provided by Lemma 5 (e.g., $|\zeta(\sigma+iT)|\ll T^{\mu(\sigma)}$)
+$\left|\int_{\sigma_{0}}^{b}f(\sigma\pm iT)\frac{x^{\sigma\pm iT}}{\sigma\pm iT}\,d\sigma\right|\ll_{r,k}\frac{x^{b}}{T}\int_{\sigma_{0}}^{b}|f(\sigma\pm iT)|\,d\sigma.$
+With $T=x$, this contribution is typically bounded by $\mathcal{O}_{r,k}\left(x^{\frac{1}{2}+\epsilon}\right)$. On the line $\text{Re}(s)=\sigma_{0}=\frac{1}{2}+\epsilon$, the integral is
 $\int_{\sigma_{0}-iT}^{\sigma_{0}+iT}f(s)\frac{x^{s}}{s}\,ds.$
-The magnitude is dominated by $x^{\sigma_{0}}=x^{\frac{1}{2}+\epsilon}$. Using the mean value theorem for the zeta-function (integral of $|\zeta(\frac{1}{2}+it)|^{2}$), we obtain
-$\int_{-T}^{T}\left|f\left(\frac{1}{2}+\epsilon+it\right)\right|\frac{x^{\frac{1}{2}+\epsilon}}{|t|}\,dt\ll x^{\frac{1}{2}+\epsilon}\cdot T^{\epsilon}\ll x^{\frac{1}{2}+\epsilon}.$
+The magnitude is dominated by $x^{\sigma_{0}}=x^{\frac{1}{2}+\epsilon}$. Using Lemma 5 and the mean value theorem for the zeta-function (integral of $|\zeta(\frac{1}{2}+it)|^{2}$), we obtain
+$\int_{-T}^{T}\left|f\left(\frac{1}{2}+\epsilon+it\right)\right|\frac{x^{\frac{1}{2}+\epsilon}}{|t|}\,dt\ll_{r,k}x^{\frac{1}{2}+\epsilon}\cdot T^{\epsilon}\ll_{r,k}x^{\frac{1}{2}+\epsilon}.$
 Combining the residue calculation and the error bounds, we obtain the equality:
-$\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=\text{Res}_{s=1}+O\left(x^{\frac{1}{2}+\epsilon}\right).$
+$\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=\text{Res}_{s=1}+\mathcal{O}_{r,k}\left(x^{\frac{1}{2}+\epsilon}\right).$
 Substituting this back into the original Perron’s formula, we rigorously establish the result of the theorem
-$\sum_{n\leq x}F_{r,k}(n)=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right).$
+$\sum_{n\leq x}F_{r,k}(n)=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+\mathcal{O}_{r,k}\left(x^{\frac{1}{2}+\epsilon}\right).$
 This completes the proof. ∎